OpenStudy (anonymous):

145/100 in lowest terms

OpenStudy (anonymous):

1 45/100 = 1 9/20

OpenStudy (anonymous):

in mixed fraction :)

OpenStudy (anonymous):

29/20

OpenStudy (anonymous):

oh just a fraction please :)

OpenStudy (anonymous):

in improper is 29/20

OpenStudy (anonymous):

not improper

OpenStudy (anonymous):

improper is like a "normal number" it's like in your question you said 145/100 = 29/20 =]

OpenStudy (anonymous):

thanks :)

OpenStudy (lalaly):

:)

OpenStudy (anonymous):

=] you welcomez :_)